Place the following equation in the standard form of a quadratic equation:
4/x^2+3/x-10=0. Please tell me what is a, b, and c.

Answer by dabanfield(803)

We need to get the equation in the standard form:
a*x^2 + b*x + c = 0
4/x^2+3/x-10=0
Multiply both sides of the equation above to clear x from the denominators:
(4/x^2)*x^2 + (3/x)*x^2 - 10*x^2 = 0*x^2
4 + 3x - 10x^2 = 0
-10x^2 + 3x + 4 = 0
a = -10
b = 3
c = 4